OpenStudy (anonymous):

Without calculating , find which is greater: 31^11 or 17^14

OpenStudy (anonymous):

@mukushla

OpenStudy (lgbasallote):

which is greater 2^4 or 3^3

OpenStudy (anonymous):

3^3

OpenStudy (lgbasallote):

so 31^11 is greater

OpenStudy (lgbasallote):

i have a 50:50 chance im right

OpenStudy (anonymous):

Base will decide who is greater ???

OpenStudy (anonymous):

*which

OpenStudy (lgbasallote):

my metaphor decided

OpenStudy (ghazi):

31^11

OpenStudy (lgbasallote):

now im depnding on probability to hope im right

OpenStudy (unklerhaukus):

unlucky Lgba

OpenStudy (lgbasallote):

because i will be seriously screwed if i wasn't

OpenStudy (anonymous):

Which is greater 2^4 or 3^2 ??

OpenStudy (ghazi):

i guess power works

OpenStudy (anonymous):

2^4

OpenStudy (anonymous):

No both base and power will work I think..

OpenStudy (lgbasallote):

i still stand by my metaphor

OpenStudy (anonymous):

I was asking lgba for his probability conclusion..

OpenStudy (lgbasallote):

i have a 50:50 chance..if i change i'll be a victim of monty hall's principle

OpenStudy (anonymous):

And what that principle says I don't know @lgbasallote

OpenStudy (ghazi):

well the answer is 17^14

OpenStudy (lgbasallote):

yes....

OpenStudy (anonymous):

The question is how @ghazi

OpenStudy (lgbasallote):

close fight

OpenStudy (ghazi):

calculator ...lol

OpenStudy (lgbasallote):

i think i now know o.O

OpenStudy (anonymous):

Without calculating you need the answer..

OpenStudy (lgbasallote):

no..nevermind

OpenStudy (lgbasallote):

how sure are we there was even any way?

OpenStudy (ghazi):

got the answer..guys ait i can do it

OpenStudy (ghazi):

*wait

OpenStudy (anonymous):

\[\frac{17^{14}}{31^{11}}>\frac{17^{14}}{34^{11}}=(\frac{1}{2})^{11}17^3>(\frac{1}{2})^{11}(2^4)^3=\frac{2^{12}}{2^{11}}=2\]

OpenStudy (anonymous):

i got it. just compared with the powers of 2

OpenStudy (ghazi):

divide both the numbers by (17)^11

OpenStudy (lgbasallote):

but you're not allowed to calculate...

OpenStudy (lgbasallote):

you can't do any math according to the instructions..you can just use critical thinking

OpenStudy (ghazi):

you'll have (2)^11 and (17)^3 so obviously (17)^3 is bigger

OpenStudy (anonymous):

Wow @mukushla great benim yoldashim...

OpenStudy (ghazi):

\[\frac{ (31)^{11} }{ (17^{11}) }<\frac{ (17)^{14} }{ (17)^{11} }\]

OpenStudy (ghazi):

\[(1.83)^{11}<(17)^{3}\]

OpenStudy (ghazi):

1.83^11=let's say 2^11=2048 and 17^3=4193...proved